OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 64.36, If the perimeter of the first polygon is 35 cm. Whats the perimeter of the second polygon? A)21.76 B)24.52 C)23.25 D)26.25

OpenStudy (anonymous):

The ratio of similar polygons: a : b The ratio of the areas of similar polygons: a^2 : b^2 \[\frac{64}{36} = \frac{a^{2}}{b^{2}}\] a : b = 8 : 6 = 4 : 3 \[\frac43 = \frac{35}{x}\] (35)(3) = 4x 105 = 4x

OpenStudy (mertsj):

If the ratio of the areas is 64 to 36, the ratio of the perimeters is 8 to 6

OpenStudy (anonymous):

Can you solve for the x?

OpenStudy (anonymous):

no i dont know how.

OpenStudy (mertsj):

So 8/6=35/x 4/3=35/x 4x=105 x=26.25

OpenStudy (anonymous):

105 = 4x 105/4 = 4x/4 26.25 = x

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

wow yes. thank you (:

OpenStudy (anonymous):

You're welcome :)